OpenStudy (chayanne):

ABCD is a rectangle. Find the length of each diagonal. AC = 2(5a + 1) and BD = 2(a + 1) I kept getting 0 and 2 but Im not sure if that right

jhonyy9 (jhonyy9):

come back please i like help you but just than collaborate - ok. ?

jhonyy9 (jhonyy9):

do you know how are the diagonales of a rectangle ? has different length ? or are equales ?

OpenStudy (chayanne):

Hello!!! ANd thank you Sorry thats all the info they give me for this equation

jhonyy9 (jhonyy9):

yes these are eligible to solve this exercise just you need to know how are the length of diagonals of a rectangle ? 

jhonyy9 (jhonyy9):

how you see there on my posted image how are these length of diagonales ? are differents ? or equales ?

jhonyy9 (jhonyy9):

do you know how many degree has every angle of a rectangle ?

jhonyy9 (jhonyy9):

angles - sorry so has 90 degree what mean that are congruent so equales - yes ?

jhonyy9 (jhonyy9):

courage

jhonyy9 (jhonyy9):

so what is your opinion please ?

OpenStudy (chayanne):

I believe they are equales. and also AB and CD are parallels too

jhonyy9 (jhonyy9):

so this is true really that AC = DB so from these given datets than you rewrite this equalitie can you solve it for ,,a" ? how many value wil get for a ?

jhonyy9 (jhonyy9):

so you know that AC = BD and you know bc. given that AC = 2(5a +1) and BD = 2(a+1) yes ?

OpenStudy (chayanne):

Yes So A = 0???

jhonyy9 (jhonyy9):

can you writing here please your calculi how you get a=0 ?

jhonyy9 (jhonyy9):

how you get above wrote zero for a ?

OpenStudy (chayanne):

Okay Hold up please. 2(5a+1)=2(a+1) 5a+1=a+1 4a=0 (Then that's were I got losted, Sorry)

jhonyy9 (jhonyy9):

exactly this is right,correct wrote and from there result that a=0 yes ?

OpenStudy (chayanne):

Yes

jhonyy9 (jhonyy9):

so now that you know the value of ,,a" so you can calculi easy the length of diagonales easy

jhonyy9 (jhonyy9):

without this easy from the end - sorry

jhonyy9 (jhonyy9):

how you calculi these length of AC and BD ?

jhonyy9 (jhonyy9):

so you know now that a=0 and you know that AC = 2(5a +1) and BD = 2(a+1) than what you need to do for get these length ?

OpenStudy (chayanne):

You have to plug the A value back into the equation???

jhonyy9 (jhonyy9):

just substitute this value of a in place of a inside this equations and will get the right length value sure

jhonyy9 (jhonyy9):

yes exactly

jhonyy9 (jhonyy9):

so in this case AC = ?

OpenStudy (chayanne):

2!?

OpenStudy (chayanne):

Thank you so much!!!

jhonyy9 (jhonyy9):

21 ? really you get 21 ? how please ? can you rewriting these your calculi how you get this 21 ?

OpenStudy (chayanne):

I mean 2

jhonyy9 (jhonyy9):

ohhh sorry AC =2 is right so than BD = ?

OpenStudy (chayanne):

Also 2

jhonyy9 (jhonyy9):

exactly - nice so do you understand it now sure clearly ?

OpenStudy (chayanne):

Yes I do. Thank you. That really helped

jhonyy9 (jhonyy9):

anytime with great pleasure just tag me bye bye

OpenStudy (chayanne):

Bye